OpenStudy (anonymous):

Use the quadratic formula to solve x^2 – 2x = 2. Can you also solve by factoring?

OpenStudy (anonymous):

no u cant u need ot use the formula because nothing can be multiplied to make -2 and add up to also equal -2...so u need to use the formula

OpenStudy (anonymous):

-2 + 1 = -2

OpenStudy (anonymous):

how can i solve this using the quadric formula? Is it (x+-2)(x-1)=0??

OpenStudy (anonymous):

quadratic formula is \[-b \pm ((\sqrt{b ^{2}-4ac})/2a)\]

OpenStudy (anonymous):

u substitue values in that equation, ur equation has to be of formAx+Bx+C=0

OpenStudy (radar):

Militarpolica, have you solved quadratics using the formula before?

OpenStudy (radar):

\[(-b \pm \sqrt{b ^{2}-4ac})/2a\]

OpenStudy (radar):

To get you started, in your quadratic equation, a=1, b=-2, and c=-2

OpenStudy (radar):

\[(2\pm \sqrt{4+8})/2=(2\pm \sqrt{12)}/2=(2\pm2\sqrt{3})/2=1\pm \sqrt{3}\]